1 Use a contour to evaluate So da 1+x4
Answer #1

Answer:- Given that Given integration is dac Itxu - let f (2) 17 z4+1 zut Put z+1=0 z z = yu (-1) z = 1/4 (cos Cos (Titan TitVSC (-top] Sf (2) d2 f (2) d2 + when Roo and f(2) is anaistic except By Jordans lemma poles inside V, Then {f(z) dz=0 Rep {3 2 راما (5) 3 (+1)3 (2) 3 (1+1) 3 دا3 = تورهای + ) - X2 2 . t) 3( ز+ ۱-)